### v4

# and thermofield double states as quantum many-body scars

Abstract
We establish a duality between open many-body systems governed by the Gorini-Kossakowski-Sudarshan-Lindblad (GKSL) equation and satisfying the detailed balance condition on the one side, and closed bilayer systems with a self-adjoint Hamiltonian on the other side. Under this duality, the identity operator on the open system side maps to a quantum many-body scar of the dual Hamiltonian $\mathcal{H}$. This scar eigenstate has a form of a thermofield double state for a single-body conserved quantity entering the detailed balance conditions. A remarkable feature of this thermofield scar is a tunable single-layer entanglement entropy controlled by the reservoir temperature on the open system side. Further, we identify broad classes of many-body open systems with nontrivial explicit eigen operators $Q$ of the Lindbladian superoperator. The expectation values of the corresponding observables exhibit a simple exponential decay, $\langle Q\rangle_{t}=e^{-\Gamma t}\langle Q\rangle_{0}$, irrespectively of the initial state. Under the above duality, these eigen operators give rise to additional (towers of) scars. Finally, we point out that more general superoperators (not necessarily of the GKSL form) can be mapped to self-adjoint Hamiltonians of bilayer systems harbouring scars, and provide an example thereof.
Introduction.  Dualities are precious gems of Physics. A duality provides a fortunate opportunity to look at the same phenomenon from different angles or, alternatively, to relate two seemingly different phenomena. Of particular value are instances when some object looks quite complex on the one side of the duality but turns out to be very simple on the other side. Then the duality becomes a powerful tool for describing this object.
Here we establish a duality between open systems described by the GKSL equation and closed systems with doubled degrees of freedom (referred to as bilayer systems throughout the paper). Dualities of this type are well known Baranger (1958); Havel (2003); Medvedyeva et al. (2016) – in a sense, their existence follows from the common mathematical nature of Lindbladian and Hamiltonian (both being linear (super-)operators) and a straightforward counting of the respective configuration space dimensions (Horn and Johnson, 1994, Chapter 4). However, in general such dualities lead to non-Hermitian Hamiltonians. In contrast, the duality presented here leads to a self-adjoint Hamiltonian, provided the open system satisfies the detailed balance condition. This way the duality connects two theories with straightforward physical meanings.
We highlight the power of the duality by relating a simple object on the open system side – the identity operator – to a nontrivial object on the closed system side – a quantum many-body scar.
Quantum many-body scars (further referred to as “scars” for brevity) are special, often explicitly known eigenstates of nonintegrable closed many-body systems Shiraishi and Mori (2017); Turner et al. (2018); Serbyn et al. (2021); Moudgalya et al. (2022). They reside in the bulk of the spectrum and typically possess distinctly nonthermal properties. Scars can manifest themselves in quantum quench experiments through the lack of thermalization and persistent oscillations of local observables Lin and Motrunich (2017); Bernien et al. (2017); Chen et al. (2022); Su et al. (2023); Zhang et al. (2023).
Remarkably, the scar obtained here turns out to be a thermofield double state with respect to a single-body observable which would be a conserved quantity in the absence of dissipation – thus we adopt the term “thermofield scar”. The thermofield double state first emerged in thermal quantum field theory Umezawa et al. (1982) and further found applications in molecular dynamics Borrelli and Gelin (2021), black hole physics Maldacena (2003); Maldacena and Susskind (2013); Hartman and Maldacena (2013) and quantum simulations Wu and Hsieh (2019); Zhu et al. (2020); Francis et al. (2021). Importantly, the single-layer bipartite entanglement entropy of the thermofield scar is thermal, with the temperature coinciding with the reservoir temperature on the open system side of the duality. In a special case of infinite temperature, the entanglement entropy is maximal and the thermofield scar reduces to the recently discovered Einstein-Podolsky-Rosen (EPR) scar Langlett et al. (2022); Wildeboer et al. (2022). At the same time, for other bipartitions the entanglement entropy of the thermofield scar is zero, which justifies the name “scar” Serbyn et al. (2021); Moudgalya et al. (2022).
Further, we present broad families of open many-body systems where nontrivial eigen operators of the Lindbladian can be explicitly constructed. The duals of these eigen operators are additional scar states of the dual bilayer system.
Finally, we discuss a more general duality that connects self-adjoint Hamiltonians of bilayer systems to superoperators that are not of the GKSL form. We provide an example where such duality enables identification of nontrivial eigenstates of the bilayer Hamiltonian.
The rest of this paper is organized as follows. We start by introducing the duality. Then we show that the duality implies the existence of the thermofield scar. After that we illustrate our general findings in a particular family of spin-$1/2$ models. Further, we demonstrate that in broad classes of dissipative models one can construct explicit eigen operators of the adjoint Lindbladian, and that these operators map into additional scars in dual bilayer closed systems. Then we briefly discuss dualities between bilayer Hamiltonians and superoperators of non-GKSL form. We conclude by discussing our results in relation to other developments and outlining further prospects of our approach.
The duality.  On the open system side, we employ the GKSL equation in the Heisenberg representation Gorini et al. (1976); Lindblad (1976),(Breuer et al., 2002, Sec. 3.2.3):
$\partial_{t}O_{t}={\mathcal{L}}^{\dagger}O_{t}\equiv i[H,O_{t}]+{\cal D}^{\dagger}O_{t}.$ (1)
Here $O_{t}$ is a Heisenberg operator of some observable, $\mathcal{L}$} is a Lindbladian superoperator (Lindbladian), $H$ is the Hamiltonian which is responsible for the coherent part of evolution, and $\cal D$ is the dissipation superoperator (dissipator) to be defined later. The operator $O_{t}$ acts in a Hilbert space $\mathscr{H}$. The expectation value $\langle O\rangle_{t}$ of the observable $O$ evolves in time according to $\langle O\rangle_{t}=\operatorname{\mbox{tr}}\rho_{0}\,O_{t}$, where $\rho_{0}$ is the initial state of the open system.
[…]
Figure 1: Illustration of the duality between an open system and a closed bilayer system. On the left an open system with a Hamiltonian $H$ and Lindblad operators $L_{j}$, $L_{j}^{\dagger}$ is schematically shown. The open system satisfies the detailed balance condition with the inverse temperature $\beta$, see eqs. (2),(3). On the right the dual closed system with doubled degrees of freedom is depicted. The Hamiltonian of the bilayer system is constructed from the Hamiltonian of the open system and the Lindblad operators, see eq. (6).
Our goal is to construct a duality that will map $i{\mathcal{L}}^{\dagger}$ to a self-adjoint Hamiltonian. To this end we employ a Wick rotation of the dissipative coupling constant on top of the linear homomorphism. We will demonstrate that one can obtain a self-adjoint Hamiltonian this way, provided the dissipator satisfies the detailed balance condition.
First we recall the general form of the GKSL dissipator, ${\mathcal{D}}^{\dagger}O_{t}=\sum_{j}\left(\tilde{L}_{j}^{\dagger}O_{t}\tilde{L}_{j}-(1/2)\{\tilde{L}_{j}^{\dagger}\tilde{L}_{j},O_{t}\}\right)$, where $\{\bullet,\bullet\}$ is an anticommutator and $\tilde{L}_{j}$ are arbitrary operators Gorini et al. (1976); Lindblad (1976); Breuer et al. (2002). The dissipator is said to satisfy the detailed balance condition if it can be written in the form (Alicki and Lendi, 2007, Sec. 1.3.4)
${\mathcal{D}}^{\dagger}O_{t}=\frac{\gamma}{2}\sum_{j}\biggl($ $e^{\frac{\beta}{2}}\left(L_{j}^{\dagger}O_{t}L_{j}-\frac{1}{2}\{L_{j}^{\dagger}L_{j},O_{t}\}\right)$
$+$ $e^{-\frac{\beta}{2}}\left(L_{j}O_{t}L_{j}^{\dagger}-\frac{1}{2}\{L_{j}L_{j}^{\dagger},O_{t}\}\right)\biggr),$ (2)
where $\beta$ is a real parameter having the meaning of the inverse reservoir temperature in microscopic derivations of the GKSL equation (Breuer et al., 2002, Sec. 3.3.2), $\gamma>0$ is the dissipative coupling constant and (rescaled) Lindblad operators $L_{j}$ are annihilation-like operators with respect to some $H_{0}$ that commutes with $H$, i.e.
$\beta H_{0},L_{j}]=-\beta L_{j},\quad[H_{0},H]=0.$ (3)
The dissipator (2) can be written in the standard GKSL form by replacing rescaled Lindblad operators $L_{j}$ by conventional Lindblad operators $\tilde{L}_{j}$ according to $\tilde{L}_{2j}=\sqrt{\gamma}e^{\frac{\beta}{4}}L_{j}$ , $\tilde{L}_{2j+1}=\sqrt{\gamma}e^{-\frac{\beta}{4}}L_{j}^{\dagger}$. Note that the detailed balance condition entails the constraint $\tilde{L}_{2j}^{\dagger}=e^{\frac{\beta}{2}}\tilde{L}_{2j+1}$.
The detailed balance condition (2), (3) naturally emerges when deriving the GKSL equation for a system interacting with a thermal reservoir in weak coupling or low density limits (Alicki and Lendi, 2007, Sec. 1.3.4). [^1] In fact, the detailed balance condition has a more general form, which can be obtained by replacing $\beta\rightarrow\beta\omega_{j}$ in the right hand sides of eqs. (2) and (3), where $\omega_{j}$ can be all different (Alicki and Lendi, 2007, Sec. 1.3.4). Further generalizations of the detailed balance in quantum systems are discussed in Fagnola and Umanita (2010). While our construction can be generalized appropriately, we leave such generalizations to further studies. Under this condition, the GKSL equation entails a classical Markov equation with classical detailed balance for populations of the conserved quantity $H_{0}$ (see e.g. (Breuer et al., 2002, Section 3.3.2)).
Note that $\beta$ is not cancelled in eq. (3) for a purpose: at infinite temperature, $\beta=0$, this equation is satisfied for arbitrary $L_{j}$. Importantly, in this case the Lindblad operators can be redefined to be self-adjoint, $L_{j}^{\prime}=(L_{j}+L_{j}^{\dagger})/\sqrt{2}$, $L_{j}^{\prime\prime}=i\,(L_{j}-L_{j}^{\dagger})/\sqrt{2}$, with the dissipator reducing to ${\mathcal{D}}^{\dagger}O_{t}=-\frac{\gamma}{2}\,\sum_{j}\big([L^{\prime}_{j},[L_{j}^{\prime},O_{t}]]+[L_{j}^{\prime\prime},[L_{j}^{\prime\prime},O_{t}]]\big).$ Vice versa, if one takes all $L_{j}$ in eq. (2) to be self-adjoint from the outset, the dissipator assumes the same form,
${\mathcal{D}}^{\dagger}O_{t}=-\frac{\gamma^{\prime}}{2}\,\sum_{j}[L_{j},[L_{j},O_{t}]],\quad L_{j}^{\dagger}=L_{j},$ (4)
where $\gamma^{\prime}=\gamma\cosh(\beta/2)$. We will refer to the dissipators of the form (4) as infinite-temperature dissipators, although they can also arise in other settings, e.g. in the singular-coupling limit (Breuer et al., 2002, Sec. 3.3.3).
To define the duality, we first fix an orthonormal basis $\{|n\rangle\}$ in the Hilbert space $\mathscr{H}$. After that we map operators $O$ on the open system side to vectors $|O\rangle\rangle_{\beta}$ on the closed system side according to
[…]
$|O\rangle\rangle_{\beta}=\sum_{m,n}O_{mn}e^{-\frac{\beta}{4}H_{0}}|m\rangle\otimes e^{-\frac{\beta}{4}H_{0}^{T}}|n\rangle,$ (5)
where the subscript $\beta$ emphasizes that the dual vector depends on the inverse temperature. This map induces the map between superoperators acting on $O$ and operators acting on $|O\rangle\rangle_{\beta}$. Applied to the Lindbladian, it reads $i\,{\mathcal{L}}^{\dagger}\,\longleftrightarrow\,\mathcal{H}_{\beta}$ with sup
$\mathcal{H}_{\beta}=$ $\left(-H-\frac{g}{4}\sum_{j}\left(e^{\frac{\beta}{2}}L_{j}^{\dagger}L_{j}+e^{-\frac{\beta}{2}}L_{j}L_{j}^{\dagger}\right)\right)\otimes\mathbb{1}$
$+\mathbb{1}\otimes\left(H^{T}-\frac{g}{4}\sum_{j}\left(e^{\frac{\beta}{2}}L_{j}^{T}L_{j}^{*}+e^{-\frac{\beta}{2}}L_{j}^{*}L_{j}^{T}\right)\right)$
$+\frac{g}{2}\sum_{j}\biggl(L_{j}^{\dagger}\otimes L_{j}^{T}+L_{j}\otimes L_{j}^{*}\biggr).$ (6)
Here the Wick rotation $g=i\gamma$ of the dissipative coupling constant has been performed. One can check directly that whenever $g$ is real, the Hamiltonian ${\cal H}_{\beta}$ is self-adjoint, as desired.
The Hamiltonian ${\cal H}_{\beta}$ is naturally interpreted as the Hamiltonian of a bilayer system: the first (the second) line of eq. (6) represents the Hamiltonian of the first (the second) layer, while the third line – the interaction between the layers, see Fig. 1 for illustration. The layers can have a direct geometrical meaning Barkeshli and Qi (2014); Hubert et al. (2019); Gall et al. (2021); Bohrdt et al. (2022) or, alternatively, represent a binary internal degree of freedom, such as fermionic spin in the Hubbard model sup .
It should be stressed that the Wick rotation along, without the similarity transformation (5), does not lead to a self-adjoint dual Hamiltonian unless all Lindblad operators can be chosen to be self-adjoint, as in eq. (4). The latter case was considered in our previous paper Teretenkov and Lychkovskiy (2024). Furthermore, a general Lindbladian can not be mapped to a self-adjoint Hamiltonian by means of any similarity transformation and/or Wick rotation. This can be seen from the fact that such mapping would imply the diagonalizability of the Lindbladian, which is not guaranteed even for a two-level system Lendi (1987). These observations highlight the crucial role of the detailed balance condition for obtaining the duality.
Note that a different choice of the basis in eq. (5) leads to a different duality. In particular, matrix transposition and complex conjugation are basis-dependent operations, therefore $H^{T}$, $L_{j}^{T}$ and $L_{j}^{*}$ entering eq. (6) depend on the basis choice. One consequence of the freedom to chose the basis is discussed in the next section.
The duality (5), (6) is the first main result of the present Letter.
The thermofield scar.  The GKSL equation in the Heisenberg picture (1) always preserves identity operator $\mathbb{1}$, i.e.
$i{\mathcal{L}}^{\dagger}\,\mathbb{1}=0.$ (7)
The duality then implies that the dual vector,
$|\mathbb{1}\rangle\rangle_{\beta}=\sum_{n}e^{-\frac{\beta}{4}H_{0}}|n\rangle\otimes e^{-\frac{\beta}{4}H_{0}^{T}}|n\rangle,$ (8)
is an (unnormalized) eigenstate of the dual Hamiltonian,
${\mathcal{H}}_{\beta}|\mathbb{1}\rangle\rangle_{\beta}=0.$ (9)
This is the second main result of the Letter.
Figure 2: Entanglement of the thermofield scar (8) with respect to different bipartitions of a bilayer system. (a) For the bipartition that singles out one layer and traces out another one, the entanglement entropy equals the thermal entanglement entropy for the Hamiltonian $H_{0}$ at temperature $\beta$. (b) The entanglement entropy vanishes for a transverse bipartition that does not cut the interlayer bonds (here the underlying basis vectors are assumed to be of product form, see the text for details).
The state (8) is a thermofield double state Umezawa et al. (1982) for $H_{0}$. It possesses a remarkable entanglement structure highly dependent on the bipartition. If one traces out one layer of the thermofield scar (8), see Fig. 2 (a), one obtains a reduced state of the second layer which turns out to be a Gibbs state with respect to $H_{0}$ (Nastase, 2023, Section 32.5), (sup, , Eq. (S8)),
$\operatorname{Tr}_{\text{one layer}}|\mathbb{1}\rangle\rangle_{\beta}\langle\langle\mathbb{1}|_{\beta}=e^{-\beta H_{0}},$ (10)
where normalization is omitted. [^2] Curiously, the reduced state (10) is a steady state of the Lindbladian $\mathcal{L}$}. As a consequence, the entanglement entropy between layers is merely the thermal equilibrium entropy for a system with the Hamiltonian $H_{0}$ and inverse temperature $\beta$.
However, for “transverse” bipartitions shown in Fig 2(b), the picture is completely different. The precise value of the entanglement entropy for such bipartitions depends on the choice of the basis underlying the duality. In the extreme case where each basis vector is of product form, the thermofield state (8) factorizes for any transverse bipartition and the entanglement entropy is equal to zero. This is explicitly demonstrated for an exemplary system considered in what follows.
In general, the double layer Hamiltonian (6) is nonintegrable and non-localized (here we assume no disorder on the open system side). Further, zero energy is, in general, in the middle of the spectrum. Therefore the zero energy eigenstate (8) possessing a zero entanglement entropy for certain bipartitions is classified as a scar Serbyn et al. (2021); Moudgalya et al. (2022). We adopt the term “thermofield scar” for this state.
We remark that in some cases, even after the thermofield scar and related towers of special eigenstates discussed below are removed from the spectrum, the rest of the spectrum may not appear completely chaotic for various reasons. In this case the term “scar” applied to the eigenstate (8) may become a misnomer, since scars are supposed to be embedded in the surrounding of chaotic eigenstates. With this caveat in mind, we nevertheless keep using the term “thermofield scar” for the sake of brevity.
A particular case of the thermofield scar (8) with $\beta=0$ has been recently discovered in refs. Langlett et al. (2022); Wildeboer et al. (2022) (without using or recognizing the duality). [^3] See also ref. Cottrell et al. (2019) where a local bilayer Hamiltonian is constructed whose ground state is approximately the thermofield double state. It has been tagged as the EPR scar in ref. Wildeboer et al. (2022), for reasons that will be clarified in the example below. Refs. Langlett et al. (2022); Wildeboer et al. (2022) contain an in-depth analysis of the various features of infinite-temperature thermofield scars, including a numerical demonstration of persistent oscillations of local observables for a quantum quench.
An example.  Consider an arbitrary-range isotropic dissipative model of $N$ spins $1/2$ with the Hamiltonian
$H=\sum_{i<j}w_{ij}\left(\sigma^{x}_{i}\sigma^{x}_{j}+\sigma^{y}_{i}\sigma^{y}_{j}\right),$ (11)
$w_{ij}$ being arbitrary real coupling constants, and the dissipator of the form (2) with
$L_{j}=\sigma^{-}_{j}\equiv(1/2)(\sigma^{x}_{j}-i\sigma^{y}_{j}).$ (12)
This model satisfies the detailed balance condition (3) with respect to the total magnetization in the $z$-direction,
$H_{0}=S^{z}=\frac{1}{2}\sum_{j}\sigma_{j}^{z},\quad[H_{0},H]=0.$ (13)
Fixing the basis to be the common eigenbasis of all $\sigma^{z}_{j}$, one obtains from eq. (6) the explicit dual self-adjoint Hamiltonian:
$\mathcal{H}_{\beta}=$ $-\sum_{i<j}w_{ij}\left(\,\sigma^{x}_{i}\sigma^{x}_{j}+\,\sigma^{y}_{i}\sigma^{y}_{j}\right)-\frac{g}{4}\sinh\frac{\beta}{2}\sum_{j}\sigma^{z}_{j}$
$+\sum_{i<j}w_{ij}\left(\,\tau^{x}_{i}\tau^{x}_{j}+\,\tau^{y}_{i}\tau^{y}_{j}\right)-\frac{g}{4}\sinh\frac{\beta}{2}\sum_{j}\tau^{z}_{j}$
$+\frac{g}{4}\sum_{j}\bigl(\sigma^{x}_{j}\tau^{x}_{j}-\sigma^{y}_{j}\tau^{y}_{j}\bigr)-N\frac{g}{2}\cosh\frac{\beta}{2}.$ (14)
Here we have changed the notations as $\sigma^{x,y,z}_{j}\otimes\mathbb{1}\rightarrow\sigma^{x,y,z}_{j}$ and $\mathbb{1}\otimes\sigma^{x,y,z}_{j}\rightarrow\tau^{x,y,z}_{j}$, so that now $\sigma^{x,y,z}_{j}$ refers to the first layer while $\tau^{x,y,z}_{j}$ – to the second one. The first (the second) line in eq. (14) is the Hamiltonian of the first (second) layer and the third line is the interlayer interaction plus an additional constant term.
The thermofield scar (8) in the considered model can be presented in the factorized form,
$|\mathbb{1}\rangle\rangle_{\beta}=\bigotimes_{j=1}^{N}\Big(e^{-\frac{\beta}{4}}|\uparrow\,\uparrow\rangle_{j}+e^{\frac{\beta}{4}}|\downarrow\,\downarrow\rangle_{j}\Big),$ (15)
where each factor is a state of two adjacent spins residing in different layers. At $\beta=0$ each factor represents a EPR pair, thus the term “EPR scar” in ref. Wildeboer et al. (2022).
Figure 3: Scars and scarred dynamics for the bilayer system (14) with $N=6$ spins in each layer. The Hamiltonian parameters are $g=1$, $\beta=0.7$, $\omega_{j\,j+1}=1$, $\omega_{j\,j+2}=0.5$ and $\omega_{ij}=0$ otherwise. (a) Rényi entropy $(-\log\operatorname{\mbox{tr}}\rho_{E}^{2})$ of eigenstates with respect to the transverse bipartition (see Fig. 2(b)) in two equal parts ($\rho_{E}$ is the corresponding reduced density matrix of an eigenstate). Highlighted are the thermofield scar $|\mathbb{1}\rangle\rangle_{\beta}$ and the scars $|Q_{n}\rangle\rangle_{\beta}$, $n=1,2,\dots,6$. (b)  Overlaps $|\langle E|\Psi_{0}\rangle|^{2}$ between the eigenstates and the product initial state $\Psi_{0}=|\downarrow\downarrow\dots\downarrow\rangle$ (all spins down in both layers). One can see that this initial state is expanded over the thermofield scar and the scars $|Q_{n}\rangle\rangle_{\beta}$. (c) Persistent oscillations of the squared total spin polarizations $\left(S^{x,y,z}\right)^{2}$ after initialization in the state $\Psi_{0}$.
Explicit eigen operators of the Lindbladian.  An adjoint Lindbladian $\mathcal{L}$$^{\dagger}$ can have other eigen operators apart from the trivial identity operator considered above. Here we focus on eigen operators $Q$ that would be integrals of motion without dissipation,
$[H,Q]=0.$ (16)
They are referred to as Hamiltonian integrals of motion (HIoM). We assume that a HIoM $Q$ is an eigen operator of the adjoint dissipator and hence of the adjoint Lindbladian,
${\mathcal{D}}^{\dagger}Q={\mathcal{L}}^{\dagger}Q=-\Gamma Q$ (17)
where $\Gamma$ is a real non-negative number. We refer to such HIoMs as eigen-HIoMs. The GKSL equation (1) is immediately solved for an eigen-HIoM $Q$ with the result
$Q_{t}=e^{-\Gamma t}Q.$ (18)
The expectation value of the eigen-HIoM experiences the same exponential decay for an arbitrary initial condition, $\langle Q\rangle_{t}=e^{-\Gamma t}\langle Q\rangle_{0}.$ Note that in the case of integrable $H$ a useful generalization of this construction exists sup .
Remarkably, there are broad classes of models where eigen-HIoMs can be explicitly found. The model (11), (12) is one example. There the condition (17) is satisfied for a set $Q_{n}$ of eigen-HIoMs that are polynomials constructed from $S^{z}$, $n\leq N$ being the polynomial degree. The first two of them read sup
$Q_{1}$ $=S^{z}+\frac{N}{2}\tanh\frac{\beta}{2}\cdot{\mathbb{1}},$ $\Gamma_{1}$ $=\gamma\cosh\frac{\beta}{2},$
$Q_{2}$ $=\left(S^{z}\right)^{2}+(N-1)\tanh\frac{\beta}{2}S^{z}$
$+\frac{N}{4}\left((N-1)\tanh^{2}\frac{\beta}{2}-1\right)\cdot{\mathbb{1}},$ $\Gamma_{2}$ $=2\,\gamma\cosh\frac{\beta}{2},$
$\dots$ (19)
In addition, the single-layer Hamiltonian $H$ is also an eigen-HIoM with $\Gamma=\gamma\cosh\frac{\beta}{2}$.
Other families of models with eigen-HIoMs are presented in the Supplement sup . They include spin and fermionic models with various dissipative terms.
Dual scars.  Explicit (towers of) eigen-HIoMs $Q$ of the adjoint Lindbladian discussed in the previous section map to (towers of) scars $|Q\rangle\rangle_{\beta}$ under the duality according to eq. (5). Remarkably, these scars can be obtained from the thermofield scar as
$|Q\rangle\rangle_{\beta}=Q(\beta)\otimes\mathbb{1}\;|\mathbb{1}\rangle\rangle_{\beta}=\mathbb{1}\otimes Q^{T}(-\beta)\;|\mathbb{1}\rangle\rangle_{\beta},$ (20)
where $Q(\beta)=e^{-\frac{\beta}{4}H_{0}}Qe^{\frac{\beta}{4}H_{0}}$, which can be verified by a direct calculation (sup, , Eq. (S7)). An infinite-temperature analog of this formula was presented in refs. Langlett et al. (2022); Wildeboer et al. (2022).
Explicit expressions for scars of the model (14) dual to eigen-HIoMs (Duality between open systems and closed bilayer systems, and thermofield double states as quantum many-body scars ) can be found in the Supplement sup . Note that they do not have a strictly vanishing entanglement entropy for the transverse bipartition. In Fig. 3 we show these scars immersed in the spectrum of the model (14). We also demonstrate that, analogously to the infinite-temperature case Langlett et al. (2022), the initial product state $\Psi_{0}=|\downarrow\downarrow\dots\downarrow\rangle$ (all spins down in both layers) has large overlaps with the thermofield scar (8) and the scars from the tower dual to HIoMs (Duality between open systems and closed bilayer systems, and thermofield double states as quantum many-body scars ), which entails persistent oscillations of the total spin projections squared, $\left(S^{\alpha}\right)^{2}(t)$, $\alpha=x,y,z$.
Curiously, the spectrum of the model (14) apparently contains other scars. Two of them, $|P_{1}\rangle=\bigotimes_{j=1}^{N}|\uparrow\,\downarrow\rangle_{j}$ and $|P_{2}\rangle=\bigotimes_{j=1}^{N}|\downarrow,\uparrow\rangle_{j}$ (spins up in one layer, spins down in another layer), are of the product form and do not depend on $\beta$. While they can be understood in terms of the duality sup , in fact they are examples of curious special eigenstates generic for a wider class of bilayer systems beyond the duality sup . We leave deeper exploration of the model (14) for further studies.
The observation that numerous families of dissipative many-body models possess (towers of) explicit isolated eigen operators of the adjoint Lindbladian, and that these eigen operators turn into (towers of) quantum many-body scars in dual bilayer systems constitute the third main result of the present paper.
Beyond GKSL.  Finally, let us briefly comment on dualities in a more general perspective. The dimension counting tells us that, in fact, an arbitrary self-adjoint bilayer Hamiltonian $\mathcal{H}$ can be mapped to some superoperator $\mathcal{S}$ acting on operators in a Hilbert space corresponding to a single layer. This superoperator is not necessarily of the GKSL form, therefore, at the first sight, such a duality misses much of its physical appeal. [^4] Note that non-Markovian dynamics of open quantum systems can lead to GKSL-like, but non-GKSL generators Hall et al. (2014). However, it can still be very useful, in particular for identifying special eigenstates of $\mathcal{H}$. To illustrate the latter claim, consider a bilayer spin-1/2 Hamiltonian $\mathcal{H}=\sum_{i<j}w_{ij}\left(\,\sigma^{x}_{i}\sigma^{z}_{j}+\tau^{x}_{i}\tau^{z}_{j}\right)+\sum_{j}\sigma^{y}_{j}\tau^{y}_{j}$ (notations are the same as in the Hamiltonian (14)). The infinite-temperature version of the map (5) leads to the superoperator ${\mathcal{S}}O=\{H,O\}-\sum_{j}\sigma^{y}_{j}\,O\,\sigma^{y}_{j}$, where $H=\sum_{i<j}w_{ij}\sigma^{x}_{i}\sigma^{z}_{j}$ and $\{\bullet,\bullet\}$ is the anticommutator. Clearly, $\mathcal{S}$} is not of the GKSL form. However, one easily identifies two eigen operators $Q_{x,z}$ of $\mathcal{S}$}, $Q_{x,z}=\prod_{j}\sigma_{j}^{x,z}$, satisfying ${\mathcal{S}}\,Q_{x,z}=N\,Q_{x,z}$. They correspond to two eigenstates $|Q_{x}\rangle\rangle=\bigotimes_{j}\left(|\uparrow\,\downarrow\rangle_{j}+|\downarrow\,\uparrow\rangle_{j}\right)$ and $|Q_{z}\rangle\rangle=\bigotimes_{j}\left(|\uparrow\,\uparrow\rangle_{j}-|\downarrow\,\downarrow\rangle_{j}\right)$ of the bilayer Hamiltonian $\mathcal{H}$ (notations are the same as in eq. (15)). This example reveals that viewing bilayer systems through the prism of a general duality of type (5) can help identify their features concealed otherwise. Exploring such dualities constitutes a promising direction for further research.
Discussion and outlook.  Eigen operators of the Lindbladian dual to scars in bilayer closed systems can be thought of as “operator scars” in open systems. [^5] See refs. Wang et al. (2024); de Leeuw et al. (2023) for somewhat different perspectives on operator scars. Their pure exponential decay is in sharp contrast to the ergodic evolution of a typical operator that explores the whole operator space allowed by symmetries Viswanath and Müller (2008); Parker et al. (2019). From a different angle, the pure exponential decay of HIoMs resembles purely oscillating behavior of decoherence free states Lidar et al. (1998); Agredo et al. (2014) and subalgebras Agredo et al. (2022). It can be interesting to look for a deeper connection between these phenomena. For the moment we just remark that the decoherence free algebras in open systems are dual to the states of closed bilayer systems with the decoupled dynamics of the layers.
A phenomenon closely related to quantum many-body scarring is Hilbert space fragmentation Žnidarič (2013); Sala et al. (2020); Khemani et al. (2020); Moudgalya et al. (2021). As we discuss in the Supplement sup , the Hilbert space fragmentation in closed bilayer systems is dual to the operator space fragmentation Essler and Piroli (2020) (see also Li et al. (2023)) in open systems. The latter phenomenon typically occurs in open systems with integrable Hamiltonians Essler and Piroli (2020); Robertson and Essler (2021); Li et al. (2023); Horstmann et al. (2013); Bakker et al. (2020); Eisler (2011); Temme et al. (2012); Shibata and Katsura (2019); Dolgirev et al. (2020); Turkeshi and Schiró (2021); Haken and Strobl (1973); Rahmani (2015); Foss-Feig et al. (2017); Budini and Garrahan (2021); de Leeuw et al. (2023); Teretenkov and Lychkovskiy (2024). Importantly, the structure of the fragmentation is much more transparent on the open system side of the duality Dodonov and Man’ko (1988); Prosen (2008); Teretenkov (2019, 2020); Nosal’ and Teretenkov (2020); Ivanov and Teretenkov (2022); Linowski et al. (2022); Nosal and Teretenkov (2022); Barthel and Zhang (2022); Zhang and Barthel (2022); Zunkovic (2014), analogously to the case of scars.
Much of the theory presented here remains valid in the case of time-dependent Hamiltonian and Lindblad operators. In particular, the thermofield state (8) remains a zero energy stationary state of the now-time-dependent dual bilayer Hamiltonian in this case, provided the detailed balance conditions (2) are satisfied at each moment of time. This paves the way to Floquet scars Mizuta et al. (2020); Yarloo et al. (2020); Mukherjee et al. (2020); Moudgalya et al. (2022) in periodically driven bilayer systems, as well as to even more exotic stationary states of bilayer Hamiltonians that depend on time aperiodically. We leave this direction of study for further work.
Much work has been made on identifying and characterizing nonequilibrium steady states of dissipative system supporting a steady current Znidaric (2010); Žnidarič (2011); Prosen (2011); Eisler (2011); Temme et al. (2012); Karevski et al. (2013); Ilievski and Zunkovic (2014); Prosen (2015); Popkov and Schütz (2017); Ribeiro and Prosen (2019). As a potentially interesting tweak, one can attempt to map such states to scars in dual bilayer systems. This might be possible, in particular, in the case of Lindblad operators that form detailed balance pairs as in ref. Trushechkin (2019).
It can be interesting to explore the freedom to use different bases for defining the duality is yet to be explored. Here we have focused on the simplest case of product state bases; more complex bases will lead to dualities with more complex properties and, in particular, to thermofield scars with more intricate entanglement structure.
Finally, we note that the reported duality might aid quantum simulations of open systems David et al. (2024) and shed new light on topological effects in open systems Niu and Wang (2024).
Acknowledgements.
Acknowledgments. We thank I. Ermakov for useful discussions. The work of O. Lychkovskiy was supported by the Foundation for the Advancement of Theoretical Physics and Mathematics “BASIS” under Grant No. 22-1-2-55-1.
Supplementary material for Letter
“Duality between open systems and closed bilayer systems,
and thermofield double states as quantum many-body scars”
by Alexander Teretenkov and Oleg Lychkovskiy
Appendix A S1. Basic formulae for dual vectors
First of all let us recall the basic formulae for the infinite-temperature case. Remark that
$|\mathbb{1}\rangle\rangle_{0}=\sum_{n}|n\rangle\otimes|n\rangle,$ (S1)
then by definition (5) at $\beta=0$ we have
$|AB\rangle\rangle_{0}=\sum_{m,n}(AB)_{mn}|m\rangle\otimes|n\rangle=\sum_{m,n}\langle m|AB|n\rangle|m\rangle\otimes|n\rangle=\sum_{m,n,k}\langle m|A|k\rangle\langle k|B|n\rangle|m\rangle\otimes|n\rangle$
$=\sum_{m,n,k}|m\rangle\langle m|A|k\rangle\otimes|n\rangle\langle n|B^{T}|k\rangle=\sum_{k}A|k\rangle\otimes B^{T}|k\rangle=A\otimes B^{T}\;|\mathbb{1}\rangle\rangle_{0}$ (S2)
for arbitrary operators $A$ and $B$. In particular, for an operator $O$ we have
$|O\rangle\rangle_{0}=O\otimes\mathbb{1}\;|\mathbb{1}\rangle\rangle_{0}=\mathbb{1}\otimes O^{T}\;|\mathbb{1}\rangle\rangle_{0}.$ (S3)
The dual vector in the finite-temperature case can be expressed in terms of one in the infinite-temperature case as
$|O\rangle\rangle_{\beta}=\sum_{m,n}O_{mn}e^{-\frac{\beta}{4}H_{0}}|m\rangle\otimes e^{-\frac{\beta}{4}H_{0}^{T}}|n\rangle=e^{-\frac{\beta}{4}H_{0}}\otimes e^{-\frac{\beta}{4}H_{0}^{T}}\;|O\rangle\rangle_{0}=e^{-\frac{\beta}{4}(H_{0}\otimes\mathbb{1}+\mathbb{1}\otimes H_{0}^{T})}|O\rangle\rangle_{0}.$ (S4)
Then using (S3) and (S2) we obtain
$|O\rangle\rangle_{\beta}=(e^{-\frac{\beta}{4}H_{0}}\otimes e^{-\frac{\beta}{4}H_{0}^{T}})(O\otimes\mathbb{1})|\mathbb{1}\rangle\rangle_{0}$ $=e^{-\frac{\beta}{4}H_{0}}O\otimes e^{-\frac{\beta}{4}H_{0}^{T}}\;|\mathbb{1}\rangle\rangle_{0}$
$=e^{-\frac{\beta}{4}H_{0}}Oe^{-\frac{\beta}{4}H_{0}}\otimes\mathbb{1}\;|\mathbb{1}\rangle\rangle_{0}=\mathbb{1}\otimes e^{-\frac{\beta}{4}H_{0}^{T}}O^{T}e^{-\frac{\beta}{4}H_{0}^{T}}\;|\mathbb{1}\rangle\rangle_{0}.$ (S5)
In particular, for $O=\mathbb{1}$ we have
$|\mathbb{1}\rangle\rangle_{\beta}=e^{-\frac{\beta}{2}H_{0}}\otimes\mathbb{1}\;|\mathbb{1}\rangle\rangle_{0}=\mathbb{1}\otimes e^{-\frac{\beta}{2}H_{0}^{T}}\;|\mathbb{1}\rangle\rangle_{0},$ (S6)
then eq. (S5) can be also written as
$|O\rangle\rangle_{\beta}=e^{-\frac{\beta}{4}H_{0}}Oe^{\frac{\beta}{4}H_{0}}\otimes\mathbb{1}\;|\mathbb{1}\rangle\rangle_{\beta}=\mathbb{1}\otimes e^{-\frac{\beta}{4}H_{0}^{T}}O^{T}e^{\frac{\beta}{4}H_{0}^{T}}\;|\mathbb{1}\rangle\rangle_{\beta}.$ (S7)
Using eq. (S1) we have
$\operatorname{Tr}_{2}|\mathbb{1}\rangle_{0}\langle\mathbb{1}|_{0}=\operatorname{Tr}_{2}\sum_{n,m}|n\rangle\langle m|\otimes|n\rangle\langle m|=\sum_{n,m}|n\rangle\langle m|\delta_{nm}=\mathbb{1},$
where $\operatorname{Tr}_{2}$ is the partial trace with respect to the second layer. Then from eq. (S6) we have
$\operatorname{Tr}_{2}|\mathbb{1}\rangle_{\beta}\langle\mathbb{1}|_{\beta}=\operatorname{Tr}_{2}\bigl((e^{-\frac{\beta}{2}H_{0}}\otimes\mathbb{1})|\mathbb{1}\rangle_{0}\langle\mathbb{1}|_{0}(e^{-\frac{\beta}{2}H_{0}}\otimes\mathbb{1})\bigr)=e^{-\frac{\beta}{2}H_{0}}\operatorname{Tr}_{2}(|\mathbb{1}\rangle_{0}\langle\mathbb{1}|_{0})e^{-\frac{\beta}{2}H_{0}}=e^{-\beta H_{0}}.$ (S8)
Similarly, the same is true if one traces out the first layer.
Appendix B S2. Proof of finite-temperature duality
From eq. (S4) and (S2) we have
$|AOB\rangle\rangle_{\beta}$ $=e^{-\frac{\beta}{4}(H_{0}\otimes\mathbb{1}+\mathbb{1}\otimes H_{0}^{T})}|AOB\rangle\rangle_{0}$
$=e^{-\frac{\beta}{4}(H_{0}\otimes\mathbb{1}+\mathbb{1}\otimes H_{0}^{T})}A\otimes B^{T}|O\rangle\rangle_{0}=e^{-\frac{\beta}{4}(H_{0}\otimes\mathbb{1}+\mathbb{1}\otimes H_{0}^{T})}A\otimes B^{T}e^{\frac{\beta}{4}(H_{0}\otimes\mathbb{1}+\mathbb{1}\otimes H_{0}^{T})}|O\rangle\rangle_{\beta}.$
for arbitrary operators $A$, $B$ and $O$. Thus, the dual Hamiltonian takes the form
$\mathcal{H}_{\beta}=e^{-\frac{\beta}{4}(H_{0}\otimes\mathbb{1}+\mathbb{1}\otimes H_{0}^{T})}\mathcal{H}e^{\frac{\beta}{4}(H_{0}\otimes\mathbb{1}+\mathbb{1}\otimes H_{0}^{T})},$ (S9)
where $\mathcal{H}$ is obtained form $i\mathcal{L}^{\dagger}$ with dissipator (2) via the linear map, which transforms $A\bullet B\rightarrow A\otimes B^{T}$ for arbitrary operators $A$ and $B$, and the Wick rotation $g=i\gamma$. Namely,
$\mathcal{H}$ $=-H\otimes\mathbb{1}+\mathbb{1}\otimes H^{T}$
$+\frac{g}{2}\sum_{j}\biggl(e^{\frac{\beta}{2}}\left(L_{j}^{\dagger}\otimes L_{j}^{T}-\frac{1}{2}L_{j}^{\dagger}L_{j}\otimes\mathbb{1}-\frac{1}{2}\mathbb{1}\otimes L_{j}^{T}L_{j}^{*}\right)+e^{-\frac{\beta}{2}}\left(L_{j}\otimes L_{j}^{*}-\frac{1}{2}L_{j}L_{j}^{\dagger}\otimes\mathbb{1}-\frac{1}{2}\mathbb{1}\otimes L_{j}^{*}L_{j}^{T}\right)\biggr).$
From eq. (3) we have
$e^{-\beta H_{0}}L_{j}e^{\beta H_{0}}=e^{\beta}L_{j},\qquad e^{-\beta H_{0}}L_{j}^{\dagger}e^{\beta H_{0}}=e^{-\beta}L_{j}^{\dagger},$
$e^{-\beta H_{0}}L_{j}L_{j}^{\dagger}e^{\beta H_{0}}=L_{j}L_{j}^{\dagger},\quad e^{-\beta H_{0}}L_{j}^{\dagger}L_{j}e^{\beta H_{0}}=L_{j}^{\dagger}L_{j}.$
Using these relations eq. (S9) takes the form of eq. (6).
Appendix C S3. Classes of dissipative models satisfying condition (17)
A dissipative model is determined by its Hamiltonian $H$ and dissipator $\mathcal{D}$}. Condition (17) is satisfied by broad classes of dissipative models with $H$ and $\mathcal{D}$} of varying complexities. We do not attempt to give an exhaustive classification of such models here; rather we provide a number of examples.
C.1 Spin $1/2$ models
Consider a system of $N$ spins $1/2$.
C.1.1 $\sigma^{z}$ dissipator
Consider $\mathcal{D}$} given by the Lindblad operators $L_{j}=\sigma^{z}_{j},~j=1,2,\dots,N$, $\beta=0$. Consider further monomials in Pauli matrices,
$V^{\alpha_{1},\dots,\alpha_{n}}_{j_{1},\dots,\,j_{n}}=\prod_{l=1}^{n}\sigma_{j_{l}}^{\alpha_{l}},\quad\alpha_{l}\in\{x,y,z\}.$ (S10)
Then
${\mathcal{D}}V^{\alpha_{1},\dots,\alpha_{n}}_{j_{1},\dots,\,j_{n}}=-2\,m\,\gamma V^{\alpha_{1},\dots,\alpha_{n}}_{j_{1},\dots,\,j_{n}},$ (S11)
where $m$ is the number of $x$- and $y$-Pauli matrices in operator (S10). A Hamiltonian that is a linear combination of operators (S10) with a fixed $m$ satisfies condition (17) with $\Gamma=2m\gamma$. In particular, Hamiltonian (11) belongs to this class (with $m=2$).
C.1.2 Isotropic dissipator
Consider an isotropic dissipator $\mathcal{D}$} given by the Lindblad operators
$L_{3j-2}=\sigma^{x}_{j},\qquad L_{3j-1}=\sigma^{y}_{j},\qquad L_{3j}=\sigma^{z}_{j},$ (S12)
where $j=1,2,\dots,N$, $\beta=0$. It is easy to verify that monomials (S10) are eigenoperators of $\mathcal{D}$},
${\mathcal{D}}V^{\alpha_{1},\dots,\alpha_{n}}_{j_{1},\dots,\,j_{n}}=-4\,n\,\gamma V^{\alpha_{1},\dots,\alpha_{n}}_{j_{1},\dots,\,j_{n}}.$ (S13)
A Hamiltonian that is a linear combination of operators (S10) with a fixed $n$ satisfies condition (17) with $\Gamma=4n\gamma$.
Let us make a comment on the relevance of the isotropic dissipator to the exactly solvable Kitaev honeycomb model Kitaev (2006). The latter model features a subspace of string-like operators of size $O(N)$ that is closed with respect to commutation with the Hamiltonian Gamayun and Lychkovskiy (2022). This space turns out to be also closed with respect to the action of isotropic dissipator (S12), as can be immediately verified. Thus, augmenting the original Kitaev model with the isotropic dissipator will lead to a dissipative model with an exactly solvable dynamics of $O(N)$ observables. Further, HIoMs of this model will satisfy the condition (17). This is an interesting direction for further research.
C.1.3 $\sigma^{\pm}$ dissipation and finite temperature
Let us consider dissipator (2) with Lindblad operators
$L_{j}=\sigma^{-}_{j}:=(1/2)(\sigma^{x}_{j}-i\sigma^{y}_{j}),\qquad j=1,2,\dots,N.$ (S14)
These Lindblad operators are neither Hermitian nor unitary, in contrast to those considered previously.
The corresponding dissipator acts on individual Pauli matrices as follows:
${\mathcal{D}}^{\dagger}\sigma^{x}_{j}$ $=-\frac{\gamma}{2}\cosh\frac{\beta}{2}\sigma^{x}_{j},\qquad{\mathcal{D}}^{\dagger}\sigma^{y}_{j}=-\frac{\gamma}{2}\cosh\frac{\beta}{2}\sigma^{y}_{j},$
${\mathcal{D}}^{\dagger}\sigma^{z}_{j}$ $=-\gamma\left(\cosh\frac{\beta}{2}\sigma^{z}_{j}+\sinh\frac{\beta}{2}\mathds{1}\right).$ (S15)
Consider the Hamiltonian (11) with an additional term describing an inhomogeneous magnetic field in $z$-direction:
$H=\sum_{i<j}\left(w_{ij}^{x}\,\sigma^{x}_{i}\sigma^{x}_{j}+w_{ij}^{y}\,\sigma^{y}_{i}\sigma^{y}_{j}\right)+\sum_{j=1}^{N}h_{j}\sigma^{z}_{j}.$ (S16)
Here $h_{j}$ are real constants. It follows from eq. (S15) that
${\mathcal{D}}^{\dagger}H=-\gamma\cosh\frac{\beta}{2}H-\gamma\sinh\frac{\beta}{2}h\cdot\mathds{1},$ (S17)
where $h=\sum_{j}h_{j}$. Due to ${\mathcal{D}}^{\dagger}\mathds{1}=0$ we have
${\mathcal{D}}^{\dagger}\left(H+\tanh\frac{\beta}{2}h\cdot\mathds{1}\right)=-\gamma\cosh\frac{\beta}{2}\left(H+\tanh\frac{\beta}{2}h\cdot\mathds{1}\right).$
Thus, condition (17) is met for
$Q=H+\tanh\frac{\beta}{2}h\cdot\mathds{1}.$
The solution of the GKSL equation in Heisenberg representation reads
$H_{t}=e^{-\gamma\cosh\frac{\beta}{2}t}\,H-\tanh\frac{\beta}{2}(1-e^{-\gamma\cosh\frac{\beta}{2}t})h\cdot\mathds{1}.$ (S18)
Now let us define
$S^{z}=\frac{1}{2}\sum_{j}\sigma^{z}_{j},$ (S19)
then from (S15) we have
${\mathcal{D}}^{\dagger}\left(S^{z}+\frac{N}{2}\tanh\frac{\beta}{2}\cdot{\mathbb{1}}\right)=-\gamma\cosh\frac{\beta}{2}\left(S^{z}+\frac{N}{2}\tanh\frac{\beta}{2}\cdot{\mathbb{1}}\right).$ (S20)
Let us calculate
${\mathcal{D}}^{\dagger}(\sigma^{z}_{j}\sigma^{z}_{k})$ $=({\mathcal{D}}^{\dagger}\sigma^{z}_{j})\sigma^{z}_{k}+\sigma^{z}_{j}({\mathcal{D}}^{\dagger}\sigma^{z}_{k})$
$=-\gamma\left(\cosh\frac{\beta}{2}\sigma^{z}_{j}+\sinh\frac{\beta}{2}\mathds{1}\right)\sigma^{z}_{k}-\gamma\sigma^{z}_{j}\left(\cosh\frac{\beta}{2}\sigma^{z}_{k}+\sinh\frac{\beta}{2}\mathds{1}\right)$
$=-2\gamma\cosh\frac{\beta}{2}\sigma^{z}_{j}\sigma^{z}_{k}-\gamma\sinh\frac{\beta}{2}\left(\sigma^{z}_{j}+\sigma^{z}_{k}\right),$
then from (S19) we have
$(S^{z})^{2}=\frac{1}{4}\sum_{j,k}\sigma^{z}_{j}\sigma^{z}_{k}=\frac{1}{4}N\mathds{1}+\frac{1}{4}\sum_{j\neq k}\sigma^{z}_{j}\sigma^{z}_{k},$
hence, we obtain
${\mathcal{D}}^{\dagger}\left((S^{z})^{2}-\frac{1}{4}N\mathds{1}\right)=-2\gamma\cosh\frac{\beta}{2}\left((S^{z})^{2}-\frac{1}{4}N\mathds{1}\right)-\gamma\sinh\frac{\beta}{2}(N-1)S^{z}.$
Taking into account eq. (S20) we have
${\mathcal{D}}^{\dagger}$ $\left(\left(S^{z}\right)^{2}+(N-1)\tanh\frac{\beta}{2}S^{z}+\frac{N}{4}\left((N-1)\tanh^{2}\frac{\beta}{2}-1\right)\cdot{\mathbb{1}}\right)$
$=2\,\gamma\cosh\frac{\beta}{2}\left(\left(S^{z}\right)^{2}+(N-1)\tanh\frac{\beta}{2}S^{z}+\frac{N}{4}\left((N-1)\tanh^{2}\frac{\beta}{2}-1\right)\cdot{\mathbb{1}}\right).$
C.1.4 $\sigma^{z}\sigma^{z}$ dissipation
There was no need to specify any particular geometric structure of interactions in the Hamiltonians considered above, since each Lindblad operator acted on a single site. Here we consider Lindblad operators acting on two neighbouring sites of some lattice, and Hamiltonians with nearest-neighbour interactions. Specifically, we consider a linear, square or cubic lattice in $D=1,2,3$ dimensions, respectively. We enumerate the bonds of the lattice by $\upsilon$ and refer to the two distinct sites corresponding to the $\upsilon$’th bond as $i_{\upsilon}$ and $j_{\upsilon}$. The dissipator is given by
$L_{\upsilon}=\sigma^{z}_{i_{\upsilon}}\sigma^{z}_{j_{\upsilon}},$ (S21)
$\beta=0$, and the Hamiltonian – by
$H=\sum_{\upsilon}\left(w_{\upsilon}^{x}\,\sigma^{x}_{i_{\upsilon}}\sigma^{x}_{j_{\upsilon}}+w_{\upsilon}^{y}\,\sigma^{y}_{i_{\upsilon}}\sigma^{y}_{j_{\upsilon}}\right).$ (S22)
$\upsilon$ runs over all bonds of the lattice in both formulae above. Note that the Hamiltonian is not necessarily translation-invariant, it is just endowed with the nearest-neighbour structure.
It is easy to see that condition (17) is satisfied here:
${\mathcal{D}}H=-4\gamma(2D-1)H.$ (S23)
It is clear that this example can be generalized to other lattices, next-to-nearest neighbour interactions, as well as Lindblad operators that act on more than two neighboring sites.
C.2 Generalized Hubbard model
Here we consider interacting fermions of two types, $b$ and $c$, on an arbitrary lattice with $N$ sites, with
$H=\sum_{i<j}\left(w_{ij}^{a}\,c^{\dagger}_{i}c_{j}+w_{ij}^{b}\,b^{\dagger}_{i}b_{j}\right)+\sum_{i,j}g_{ij}\,c^{\dagger}_{i}c_{i}\,b^{\dagger}_{j}b_{j}$ (S24)
and dissipator (2)
$L_{2j-1}=b_{j},\qquad L_{2j}=c^{\dagger}_{j}c_{j},\qquad j=1,2,\ldots,N.$ (S25)
The Hamiltonian (S24) generalizes the Hubbard model.
It is easy to see that
${\mathcal{D}}\left(n_{b}+\frac{N}{e^{\beta}-1}\right)$ $=-\gamma\sinh\frac{\beta}{2}\left(n_{b}+\frac{N}{e^{\beta}-1}\right),$
${\mathcal{D}}(n_{c})$ $=0,$ (S26)
where $n_{b}=\sum_{j}b_{j}^{\dagger}b_{j}$ and $n_{c}=\sum_{j}c_{j}^{\dagger}c_{j}$.
Eq. (C.2) entails that condition (17) is satisfied for $Q_{b}=n_{b}+\frac{N}{e^{\beta}-1}$ with $\Gamma_{b}=\gamma\sinh\frac{\beta}{2}$ and $Q_{c}=n_{c}$ with $\Gamma_{c}=0$.
Such a dissipator is in detailed balance with $H_{0}=n_{a}$. Remark that it is not hard to generalize this model to the case where $L_{2N+j}=c_{j}$ are also included in dissipator (2), and, moreover, they can be included at a different temperature. Then a multitemperature generalization of (5) with $\beta H_{0}\rightarrow\beta_{a}n_{a}+\beta_{b}n_{b}$ should be done here to obtain the Hermitian Hamiltonian in the dual picture.
Appendix D S.4 Generalization of eq. (18)
Here we present a useful generalization of conditions (16),(17) and eq. (18). Consider an observable $O$ with the following property: the Heisenberg operator $O_{t}|_{\gamma=0}$ in the absence of dissipation satisfies
${\mathcal{D}}^{\dagger}\left(O_{t}|_{\gamma=0}\right)=-\Gamma\,\,O_{t}|_{\gamma=0}.$ (S27)
One immediately verifies that in this case
$O_{t}=e^{-\Gamma t}\,\left(O_{t}|_{\gamma=0}\right)$ (S28)
is the solution of the GKSL equation (1) with a finite dissipation. This generalization is important when $\left(O_{t}|_{\gamma=0}\right)$ is known exactly, e.g. as in Refs. Prosen (2000); Lychkovskiy (2021); Gamayun and Lychkovskiy (2022). In Ref. Teretenkov and Lychkovskiy (2024) eq. (S28) was applied to a dissipative $XX$ model. We remark that a similar exponential damping on top of a coherent dynamics has been found theoretically van Caspel and Gritsev (2018); Maruyoshi et al. (2023) and experimentally Maruyoshi et al. (2023) in finite $XXZ$ spin chains with dissipation.
Appendix E S.5 Operator space fragmentation
For several models of open quantum systems Dodonov and Man’ko (1988); Prosen (2008); Zunkovic (2014); Teretenkov (2019, 2020); Nosal’ and Teretenkov (2020); Ivanov and Teretenkov (2022); Linowski et al. (2022); Nosal and Teretenkov (2022); Barthel and Zhang (2022); Zhang and Barthel (2022); Teretenkov and Lychkovskiy (2024) it was shown that for some operators $R^{n}$
${\mathcal{L}}^{\dagger}R^{n}=\sum_{m}\mathcal{M}_{nm}R^{m},$ (S29)
where $\mathcal{M}_{nm}$ are numbers and the number of such operators rises polynomially with the system size  [^6] If one takes as $R_{n}$ a full matrix basis for a many-body system with finite-dimensional Hilbert space, then $\mathcal{M}_{nm}$ will be just $\mathcal{L}$$^{\dagger}$ written in such a basis, but the number of elements of such a basis rises exponentially in the system size.. If we combine the operators $R^{n}$ in one operator-valued vector $R$, then eq. (S29) takes the form
${\mathcal{L}}^{\dagger}(R)=\mathcal{M}R.$ (S30)
Moreover, for above-mentioned models the products of operators $R_{n}$ also satisfy an equation similar to eq. (S29). If one combines such products in tensors $\otimes^{k}R$, then it can be expressed
${\mathcal{L}}^{\dagger}(\otimes^{k}R)=\mathcal{M}^{(k)}(\otimes^{k}R),$ (S31)
where $\mathcal{M}^{(k)}$ transforms tensors $(\otimes^{k}R)$ to tensors of the same size. For unitary dynamics eq. (S31) follows directly from eq. (S29) due to the Leibniz rule for a commutator, which gives $\mathcal{M}^{(k)}=\sum_{j=1}^{k}\mathcal{M}_{j}$, where $\mathcal{M}_{j}$ acts as $\mathcal{M}$ in the $j$-th subspace and as identity in other tensor multiplicands. As it holds for unitary dynamics one can easily see that eq. (S31) holds for the Lindbladian with unitary Lindblad operators as well Teretenkov (2020); Nosal’ and Teretenkov (2020); Linowski et al. (2022). In case of Hermitian Linblad operators GKSL dissipators reduce to the sum of double commutators, which also led to eq. S31 due to the Leibniz rule Ivanov and Teretenkov (2022). In Ref. Teretenkov and Lychkovskiy (2024) Leibniz-like identity for the commutator of anticommutator led to the symmetrized analog of eq. (S31) for wide classes of dissipative $XX$ models and Onsager strings as elements of $R$. In Ref. Nosal and Teretenkov (2022) a generalized Leibniz-like rule for Lindbladians led to a hierarchical analog of eq. (S31), where ${\mathcal{L}}^{\dagger}(\otimes^{k}R)$ is a span of $\otimes^{j}R$ for $j=1,\ldots,k$ rather than just $\otimes^{k}R$.
If the number of operators $R_{n}$ rises polynomially in the system size, then the number of $\otimes^{k}R$ components also does. If $\otimes^{k}R$ are independent for a large enough system size, then operator space contains exponentially (in the system size) many blocks with closed dynamics, which constitutes operator space fragmentation Essler and Piroli (2020). Duality (5) brings such operator spaces into Hilbert spaces
$\operatorname{span}|\otimes^{k}R\rangle\rangle_{\beta}:=\operatorname{span}|R^{n_{1}}\cdot\ldots\cdot R^{n_{k}}\rangle\rangle_{\beta}$ (S32)
invariant under dynamics with dual Hamiltonian (6) inducing Hilbert space fragmentation for such a Hamiltonian. Let us emphasize that the structure of fragmented operator spaces induced by operators products $R^{n_{1}}\cdot\ldots\cdot R^{n_{k}}$ or symmetrized products is more evident than the structure of the dual fragmented Hilbert spaces.
Appendix F S.6 First few states in the tower of scars
Here we provide explicit expressions for two scars of the model (14) dual to eigen-HIOMS (Duality between open systems and closed bilayer systems, and thermofield double states as quantum many-body scars ). According to eq. (20), these (unnormalized) scar eigenstates and eigenenergies read
$|Q_{1}\rangle\rangle_{\beta}$ $=|S^{z}\rangle\rangle_{\beta}+\frac{N}{2}\tanh\frac{\beta}{2}\,|{\mathbb{1}}\rangle\rangle_{\beta},$ $E_{1}$ $=-g\cosh\frac{\beta}{2},$
$|Q_{2}\rangle\rangle_{\beta}$ $=|\left(S^{z}\right)^{2}\rangle\rangle_{\beta}+(N-1)\tanh\frac{\beta}{2}|S^{z}\rangle\rangle_{\beta}+\frac{N}{4}\left((N-1)\tanh^{2}\frac{\beta}{2}-1\right)\,|{\mathbb{1}}\rangle\rangle_{\beta},$ $E_{2}$ $=-2\,g\cosh\frac{\beta}{2},$ (S33)
with $|{\mathbb{1}}\rangle\rangle_{\beta}$ given by eq. (8) and
$|S^{z}\rangle\rangle_{\beta}$ $=\frac{1}{2}\sum_{j=1}^{N}\left(\Big(e^{-\frac{\beta}{4}}|\uparrow\,\uparrow\rangle_{j}-e^{\frac{\beta}{4}}|\downarrow\,\downarrow\rangle_{j}\Big)\bigotimes_{\begin{subarray}{c}l=1\\ l\neq j\end{subarray}}^{N}\Big(e^{-\frac{\beta}{4}}|\uparrow\,\uparrow\rangle_{l}+e^{\frac{\beta}{4}}|\downarrow\,\downarrow\rangle_{l}\Big)\right),$
$|\left(S^{z}\right)^{2}\rangle\rangle_{\beta}$ $=\frac{N}{4}\,|{\mathbb{1}}\rangle\rangle_{\beta}+\frac{1}{4}\sum_{i<j}\left(\Big(e^{-\frac{\beta}{4}}|\uparrow\,\uparrow\rangle_{i}-e^{\frac{\beta}{4}}|\downarrow\,\downarrow\rangle_{i}\Big)\otimes\Big(e^{-\frac{\beta}{4}}|\uparrow\,\uparrow\rangle_{j}-e^{\frac{\beta}{4}}|\downarrow\,\downarrow\rangle_{j}\Big)\bigotimes_{\begin{subarray}{c}l=1\\ l\neq i,j\end{subarray}}^{N}\Big(e^{-\frac{\beta}{4}}|\uparrow\,\uparrow\rangle_{l}+e^{\frac{\beta}{4}}|\downarrow\,\downarrow\rangle_{l}\Big)\right).$ (S34)
As noted in the main text, an additional scar of the model (14) can be obtained from the single-layer Hamiltonian $H$. This scar explicitely reads
$|H\rangle\rangle_{\beta}$ $=2\sum_{i<j}w_{ij}\left(\Big(|\uparrow\,\downarrow\rangle_{i}\otimes|\downarrow\,\uparrow\rangle_{j}+|\downarrow\,\uparrow\rangle_{i}\otimes|\uparrow\,\downarrow\rangle_{j}\Big)\bigotimes_{\begin{subarray}{c}l=1\\ l\neq i,j\end{subarray}}^{N}\Big(e^{-\frac{\beta}{4}}|\uparrow\,\uparrow\rangle_{l}+e^{\frac{\beta}{4}}|\downarrow\,\downarrow\rangle_{l}\Big)\right).$ (S35)
Note that these scars have nonzero entanglement for the transverse bipartition, in contrast to the thermofield scar $|{\mathbb{1}}\rangle\rangle_{\beta}$.
Appendix G S.7 A different type of scar states in bilayer systems
Consider a bilayer system with the Hamiltonian
$\mathcal{H}=$ $H_{1}+H_{2}+\sum_{i=1}^{N_{1}}\sum_{j=1}^{N_{2}}g_{ij}\bigl(\sigma^{x}_{i}\tau^{x}_{j}-\sigma^{y}_{i}\tau^{y}_{j}\bigr),$ (S36)
where $H_{1}$ ($H_{2}$) is a Hamiltonian of $N_{1}$ ($N_{2}$) spins $1/2$ constituting the first (second) layer, and the third term in $\mathcal{H}$ is the interlayer interaction. The Hamiltonians $H_{1}$ and $H_{2}$ conserve the $z$-projections $S^{z}_{1}=\sum_{i=1}^{N_{1}}\sigma^{z}_{i}$ and $S^{z}_{2}=\sum_{j=1}^{N_{2}}\sigma^{z}_{j}$ of spins of respective layers. Apart from this, $H_{1}$ and $H_{2}$ are arbitrary (in particular, can contain a different number of spins), and therefore the closed system $\mathcal{H}$ is not covered by the duality discussed in the main text. Still, the states
$|P_{1}\rangle=|\underbrace{\uparrow\uparrow\dots\uparrow}_{1\mathrm{st~layer}}\rangle\otimes|\underbrace{\downarrow\downarrow\dots\downarrow}_{2\mathrm{nd~layer}}\rangle,\qquad|P_{2}\rangle=|\underbrace{\downarrow\downarrow\dots\downarrow}_{1\mathrm{st~layer}}\rangle\otimes|\underbrace{\uparrow\uparrow\dots\uparrow}_{2\mathrm{nd~layer}}\rangle$ (S37)
are eigenstates of $\mathcal{H}$. Indeed, each of this states is a maximal(minimal)-eigenvalue eigenstate of $S^{z}_{1}$ and $S^{z}_{2}$ and hence of $H_{1}$ and $H_{2}$. Further, each of this states is a zero-eigenvalue eigenstate of each term of the form $(\sigma^{x}_{i}\tau^{x}_{j}-\sigma^{y}_{i}\tau^{y}_{j})$, which completes the proof.
In a special case of $\mathcal{H}$ dual to an open system according to eq. (6), the eigenstates $|P_{1}\rangle$ and $|P_{2}\rangle$ are dual to non-Hermitian eigen operators of the Lindbladian
$P_{1}=\prod_{j=1}^{N}\sigma_{j}^{+},\qquad P_{2}=\prod_{j=1}^{N}\sigma_{j}^{-}.$ (S38)